OpenStudy (anonymous):

A company produces doll houses and sets of doll furniture. The doll houses take 3 hours of labor to produce, and the furniture sets take 8 hours. The labor available is limited to 400 hours per week, and the total production capacity is 100 items per week. Existing orders require that at least 20 doll houses and 10 sets of furniture be produced per week. Write a system of inequalities representing this situation, where x is the number of doll houses and y is the number of furniture sets.

OpenStudy (jackellyn):

x = number of doll houses; y = furniture sets : The labor constraint: 3x + 8y =< 400 Arrange in the general (y=) form so we can plot the graph 8y =< 400 - 3x y =< (400/8) - (3/8)x y =< 50 - (3/8)x : The production constraint: x + y =< 100 y =< 100 - x : Min house constraint: x => 20 : Min furniture constraint y => 10